OpenStudy (anonymous):

The temperature of a body is measured as 104oF. It is observed that the amount the temperature changes for each period of two hours is -0.3 time the difference between the previous period's temperature and the room temperature, which is 70oF. a. Write a recurrence relation for tn, the temperature of the body at the end of n 2-hour time periods.

OpenStudy (anonymous):

dT/dt = -0.3(T - 70)

OpenStudy (anonymous):

\[t_{0} = 104\]

OpenStudy (anonymous):

I had T(n) = -.03(104 - 70)

OpenStudy (anonymous):

so same thing

OpenStudy (anonymous):

u know differential eqns?

OpenStudy (anonymous):

oooo...somewhat...has been a while

OpenStudy (anonymous):

dT/(T-70) = -0.3dt

OpenStudy (anonymous):

\[\int\limits_{104}^{T}dT/(T-70) = -0.3\int\limits_{0}^{t}dt\]

OpenStudy (anonymous):

need to keep it basice with recurrence relations

OpenStudy (anonymous):

ln (t-70) = -0.3ln(34)t

OpenStudy (anonymous):

thanks though!

OpenStudy (amistre64):

so initial t{0} = 140 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

t{n} = -0.3*[ t{n-1} - 70]

OpenStudy (anonymous):

I havee that but can't get to work on calc

OpenStudy (amistre64):

i havent used the calulator :) plug it in google to create a table if need be

OpenStudy (anonymous):

gues i 'll do it by hand

OpenStudy (anonymous):

how on google...tried wolfram once...pretty cool

OpenStudy (anonymous):

Find a general and particular solution for the system and give the value of t(12), the temperature of the body after 24 hours.

OpenStudy (amistre64):

i just reiterete it over and over in the right place on google

OpenStudy (amistre64):

-.3*( 104 - 70) -.3(-.3*( 104 - 70)-70) -.3(-.3(-.3*( 104 - 70)-70)-70) like that